# Legitimizing, Developing, and Sustaining Feminist HCI in East Asia: Challenges and Opportunities

### v1

@@ 1. Background and Motivation
Feminist HCI, coined by Shaowen Bardzell (bardzell2010a; bardzell2011b) in 2010, advocates for applying feminist theories and values to the design, critique, and evaluation of interactive technologies. For over a decade, this agenda has inspired a rich spectrum of studies, including the design of tools that empower women (pcos_women_health; feminism_care; sun2025; geng2025; park2025; henriques2025), LGBTQ+ communities (cui2022; taylor2024; fiesler2016archive), and other marginalized groups (ibrahim2024d; ibrahim2024c; oguamanam2023; petterson2024), as well as the critique of technologies and interfaces (grimme2024; tuli2022; pcos_women_health; india_menstruation_taboo). In the wake of this influential agenda, a growing number of scholars with East Asian backgrounds are expanding Feminist HCI research. Prior studies, for example, have examined localized forms of digital feminism practices (deng2024a; qin2024; wan2025a; pan2025; steen2023you), women’s health and well-being as shaped by regional customs and taboos (sun2025; geng2025; mandai2025; wang2024j), or equality of labor and work under regional economic and gendered norms (tang2022; zhao2024a; zhou2023). These works contributed valuable, situated knowledge to the global community by highlighting the historical, political, and cultural dynamics unique to East Asia.
@@
However, the very factors that ground these unique perspectives also create significant survival challenges for researchers in the region. Driven by pragmatism and utilitarian priorities (shen2005), funding agencies often sideline Feminist HCI research, perceiving it as less “technical” or “cutting-edge” and less valuable than productivity-oriented technologies. At the same time, feminism is frequently framed as politically sensitive in East Asia (niranjana2007feminism; heng2013great), and related scholarship such as Feminist HCI is often stigmatized, met with backlash, and at times even deemed “inappropriate” for serious funding (choi2024a). The persistent marginalization has significantly constrained the field’s growth and legitimacy. Consequently, local senior Feminist HCI researchers and established communities remain scarce in East Asia, leaving junior scholars with limited guidance and few clear career paths to follow.
@@
Yet, even in the face of these challenges, we are witnessing a vibrant surge of interest in Feminist HCI within East Asian contexts (wan2025a; deng2024a; qin2024; sun2025; tang2022; mandai2025). In addition to the increasing number of relevant publications, this momentum has also been reflected in the organizers’ community-building practices. At CHI 2025, for example, several junior co-organizers first connected through shared interests. Through personal networks, we took the initiative to launch an informal Feminist HCI Interest Group among Asian researchers. This group attracted around 30 participants to a Feminist HCI lunch gathering in Yokohama, Japan. The initiative has since evolved into an ongoing Feminist HCI Reading Seminar Group, which now includes 102 members.
@@ 3. Organizers
Runhua Zhang is a Ph.D. student at The Hong Kong University of Science and Technology. Her research in Feminist HCI focuses on gender bias in AI, digital feminism (pan2025), and women’s health (sun2025). As an advocate, she initiated informal Feminist HCI meet-ups at CHI 2025 and currently organizes a Feminist HCI Reading Seminar with 102 members.
Ruyuan Wan is a Ph.D. student at The Pennsylvania State University. Her work examines digital feminism on social media. Her CHI 2025 paper on women’s hashtag re-appropriation for audience control on Rednote received a Best Paper Honorable Mention (wan2025a). She also co-organized the informal Queer in AI meet-ups at CHI 2024 and 2025.
@@
Daye Kang is a Ph.D. candidate at Cornell University. She explores Human-AI collaboration and women’s health. Her CHI 2025 paper on tracking and managing PCOS (polycystic ovary syndrome) received the Best Paper Award (pcos_women_health).
Yigang Qin is a Ph.D. student at Syracuse University. He studies how digital technologies configure labor and work, and his CSCW 2024 paper on gender-blindness and feminist activism received the DEI Recognition Award (qin2024).
Yijia Wang is a Ph.D. student at Institute of Science Tokyo. Her research examining how East Asian cultural concepts such as kawaii (Japanese “cute”) shape gender bias or stereotypes in technologies (wang2024j).
Ziqi Pan is a Ph.D. candidate at The Hong Kong University of Science and Technology. She studies human-human interaction through daily technologies. Her work explores the well-being of queer and women’s online communities (pan2025).
@@ 4. Community of Interest
While this meet-up topics are primarily focused on Feminist HCI research in East Asian contexts, it is open to all CHI 2026 attendees by nature. Although research in this area has been steadily increasing, the visibility of East Asian scholars and research within the global CHI community remains limited (gamage2025; lu2024b). This event provides an important opportunity to bring together both scholars from the region and those with broader interests. By fostering discussion on unique challenges such as funding constraints and regional socio-political dynamics, our meet-up seeks to generate and share knowledge on how these challenges can be addressed from both regional and global perspectives. We believe such dialogues align closely with and will contribute to CHI’s core values such as diversity and inclusivity.

### v3

@@ 1. Background and Motivation
Feminist HCI, coined by Shaowen Bardzell (Bardzell, 2010; Bardzell and Bardzell, 2011) in 2010, advocates for applying feminist theories and values to the design, critique, and evaluation of interactive technologies. For over a decade, this agenda has inspired a rich spectrum of studies, including the design of tools that empower women (Kang et al., 2025; Park, 2024; Sun et al., 2025; Geng et al., 2025; Park et al., 2025; Henriques et al., 2025), LGBTQ+ communities (Cui et al., 2022; Taylor et al., 2024; Fiesler et al., 2016), and other marginalized groups (Ibrahim et al., 2024; Ibrahim, 2024; Oguamanam et al., 2023; Petterson et al., 2024), as well as the critique of technologies and interfaces (Grimme et al., 2024; Tuli et al., 2022; Kang et al., 2025; Tuli et al., 2019). In the wake of this influential agenda, a growing number of scholars with East Asian backgrounds are expanding Feminist HCI research. Prior studies, for example, have examined localized forms of digital feminism practices (Deng et al., 2024; Qin et al., 2024; Wan et al., 2025; Pan et al., 2025; Steen et al., 2023), women’s health and well-being as shaped by regional customs and taboos (Sun et al., 2025; Geng et al., 2025; Mandai et al., 2025; Wang and Seaborn, 2024), or equality of labor and work under regional economic and gendered norms (Tang et al., 2022; Zhao, 2024; Zhou et al., 2023). These works contributed valuable, situated knowledge to the global community by highlighting the historical, political, and cultural dynamics unique to East Asia.
@@
However, the very factors that ground these unique perspectives also create significant survival challenges for researchers in the region. Driven by pragmatism and utilitarian priorities (Shen and Williams, 2005), funding agencies often sideline Feminist HCI research, perceiving it as less “technical” or “cutting-edge” and less valuable than productivity-oriented technologies. At the same time, feminism is frequently framed as politically sensitive in East Asia (Niranjana, 2007; Heng, 1997), and related scholarship such as Feminist HCI is often stigmatized, met with backlash, and at times even deemed “inappropriate” for serious funding (Choi et al., 2024). The persistent marginalization has significantly constrained the field’s growth and legitimacy. Consequently, local senior Feminist HCI researchers and established communities remain scarce in East Asia, leaving junior scholars with limited guidance and few clear career paths to follow.
@@
Yet, even in the face of these challenges, we are witnessing a vibrant surge of interest in Feminist HCI within East Asian contexts (Wan et al., 2025; Deng et al., 2024; Qin et al., 2024; Sun et al., 2025; Tang et al., 2022; Mandai et al., 2025). In addition to the increasing number of relevant publications, this momentum has also been reflected in the organizers’ community-building practices. At CHI 2025, for example, several junior co-organizers first connected through shared interests. Through personal networks, we took the initiative to launch an informal Feminist HCI Interest Group among Asian researchers. This group attracted around 30 participants to a Feminist HCI lunch gathering in Yokohama, Japan. The initiative has since evolved into an ongoing Feminist HCI Reading Seminar Group, which now includes 102 members.
@@ 3. Organizers
Runhua Zhang is a Ph.D. student at The Hong Kong University of Science and Technology. Her research in Feminist HCI focuses on gender bias in AI, digital feminism (Pan et al., 2025), and women’s health (Sun et al., 2025). As an advocate, she initiated informal Feminist HCI meet-ups at CHI 2025 and currently organizes a Feminist HCI Reading Seminar with 102 members.
Ruyuan Wan is a Ph.D. student at The Pennsylvania State University. Her work examines digital feminism on social media. Her CHI 2025 paper on women’s hashtag re-appropriation for audience control on Rednote received a Best Paper Honorable Mention (Wan et al., 2025). She also co-organized the informal Queer in AI meet-ups at CHI 2024 and 2025.
@@
Daye Kang is a Ph.D. candidate at Cornell University. She explores Human-AI collaboration and women’s health. Her CHI 2025 paper on tracking and managing PCOS (polycystic ovary syndrome) received the Best Paper Award (Kang et al., 2025).
Yigang Qin is a Ph.D. student at Syracuse University. He studies how digital technologies configure labor and work, and his CSCW 2024 paper on gender-blindness and feminist activism received the DEI Recognition Award (Qin et al., 2024).
Yijia Wang is a Ph.D. student at Institute of Science Tokyo. Her research examining how East Asian cultural concepts such as kawaii (Japanese “cute”) shape gender bias or stereotypes in technologies (Wang and Seaborn, 2024).
Ziqi Pan is a Ph.D. candidate at The Hong Kong University of Science and Technology. She studies human-human interaction through daily technologies. Her work explores the well-being of queer and women’s online communities (Pan et al., 2025).
@@ 4. Community of Interest
While this meet-up topics are primarily focused on Feminist HCI research in East Asian contexts, it is open to all CHI 2026 attendees by nature. Although research in this area has been steadily increasing, the visibility of East Asian scholars and research within the global CHI community remains limited (Gamage et al., 2025; Lu et al., 2024). This event provides an important opportunity to bring together both scholars from the region and those with broader interests. By fostering discussion on unique challenges such as funding constraints and regional socio-political dynamics, our meet-up seeks to generate and share knowledge on how these challenges can be addressed from both regional and global perspectives. We believe such dialogues align closely with and will contribute to CHI’s core values such as diversity and inclusivity.